### v1

# Uncertainty Quantification for Evaluating Machine Translation Bias

Abstract
In machine translation (MT), when the source sentence includes a lexeme whose gender is not overtly marked, but whose target-language equivalent requires gender specification, the model must infer the appropriate gender from the context and/or external knowledge. Studies have shown that MT models exhibit biased behaviour, relying on stereotypes even when they clash with contextual information. We posit that apart from confidently translating using the correct gender when it is evident from the input, models should also maintain uncertainty about the gender when it is ambiguous. Using recently proposed metrics of semantic uncertainty, we find that models with high translation and gender accuracy on unambiguous instances do not necessarily exhibit the expected level of uncertainty in ambiguous ones. Similarly, debiasing has independent effects on ambiguous and unambiguous translation instances. [^1] The code will be available at https://anonymous.4open.science/r/uncertainty_bias_ambiguity-8A4C/
Uncertainty Quantification for Evaluating Machine Translation Bias
Ieva Raminta Staliūnaitė University of Cambridge irs38@cam.ac.uk                        Julius Cheng University of Cambridge jncc3@cam.ac.uk                        Andreas Vlachos University of Cambridge av308@cam.ac.uk
1 Introduction
Figure 1: Probabilities for feminine and masculine determiners in a Spanish translation of a sentence containing a noun that is either feminine (referred to as ’she’) or ambiguous (’they’), by two existing models and the ideal expected attribution of an unbiased model.
Language is inherently ambiguous, and meaning is often resolved through context. However, not all ambiguity is resolvable (van Deemter, 1998). When humans process language, they draw on linguistic, cognitive, and social biases to arrive at an interpretation (Cairns, 1973). While linguistic biases ease cognitive processing, some can also have harmful effects, such as reinforcing existing social inequalities (Beukeboom, 2013). NLP models exhibit sensitivity to many of the human biases (Echterhoff et al., 2024), and even exaggerate them (Dhamala et al., 2021), as well as introduce additional ones (Tjuatja et al., 2024). When the input is unresolvably ambiguous, favoring a single output necessarily relies on biases. Therefore, a well-designed model should refrain from making a single prediction, instead requesting clarification or generating multiple alternative outputs.
Most studies on decoding with language models (LMs) for machine translation (MT) evaluate a single prediction, usually generated with beam search, per translation instance. Consequently, most studies on uncertainty quantification (UQ) use uncertainty to predict the quality of the translation recovered by beam search (Fomicheva et al., 2020; Cheng and Vlachos, 2024). Previous work on bias in MT has focused on LM performance against gold standard labels (Stanovsky et al., 2019), and previous work on ambiguity in MT has likewise focused on resolvable cases (Barua et al., 2024; Martelli et al., 2025). Thus, most work on uncertainty and ambiguity assumes that there is a single correct translation per instance. However, less attention has been given to ambiguous source sentences, where the choice of the LM cannot be guaranteed to be correct without additional context. This may be seen as a form of aleatoric uncertainty (Hora, 1996), uncertainty which is inherent in the data and irreducible. According to Baan et al. (2024), the spread of probability mass in LMs represents both lack of confidence as well as variation in human generations. We adopt this line of thinking and examine whether LMs accurately and fairly represent the range of possible translations for ambiguous source sentences.
In this work, we leverage distribution-level uncertainty metrics to evaluate cases where the model should not be certain about its predictions due to ambiguity in the input. Figure 1 shows two versions of a sentence, with unambiguous (top) or ambiguous (bottom) gender of the noun ‘protester’, and the different probabilities assigned to the Spanish translations of the determiner of this noun. An ideal model should assign a higher probability for the feminine determiner (‘la’) when the gender is disambiguated by the pronoun, and produce equal probabilities for masculine and feminine translations when the gender is ambiguous. However, state-of-the-art MT models, including debiased ones, tend to produce more uniform probability distributions for unambiguous inputs, and less uniform distributions for ambiguous ones. This indicates that the model probabilities are influenced by stereotypical associations between protesting and masculinity, causing the models to default to the masculine form even when no gender preference is warranted, and to select the feminine form with low confidence despite clear contextual cues.
To study this systematically, we focus on translating sentences from a language that does not mark gender in nouns and verbs (English) into languages that do (Spanish, French, Ukrainian, and Russian). We use the WinoMT dataset (Stanovsky et al., 2019), which includes stereotypical gender roles, and extend it with manual translations and automatic annotations of additional cognitive bias cues, such as implicit causality verbs. In some cases, the gender is resolvable from context, while in others it is not. We explore different variants of Semantic Uncertainty metrics (Cheng and Vlachos, 2024; Farquhar et al., 2024) to quantify the semantic diversity of translation samples, finding that these metrics effectively capture the variation in gender caused by bias triggers. We validate the metrics against the established gender accuracy metric and account for the effects of translation accuracy. Our main findings are: 1) stereotypes and linguistic biases influence gender translation, 2) the degree of bias corresponds to overall model translation accuracy in unambiguous cases, 3) the degree of bias corresponds to translation accuracy at the instance level in ambiguous cases, 4) debiasing effects vary depending on input ambiguity, translation accuracy, and target language.
2 Related Work
Researchers have addressed various biases in MT, including algorithmic bias (Vanmassenhove et al., 2021), as well as gender, number and formality biases (Měchura, 2022). Gender stereotypes are triggered not only by semantic content, but also by speech mannerisms (Dawkins et al., 2024) and person names (Saunders and Olsen, 2023). Existing solutions for gender translation in both unambiguous (Robinson et al., 2024) and ambiguous cases (Cho et al., 2019; Gonen and Webster, 2020; Vanmassenhove and Monti, 2021) rely on tools or human annotation of gender, limiting their generalisability to other types of ambiguities. For example, Cho et al. (2019) explore unresolvable ambiguity by generating multiple translations for a sentence with an ambiguous pronoun, however, it is restricted to specific sentence types.
In NLP and ML research, methods for distinguishing aleatoric and epistemic uncertainty have been proposed (Hou et al., 2024), however they do not distinguish between data randomness and data ambiguity. Some work has made the link between biases and uncertainty (Sicilia et al., 2024; Kuzucu et al., 2025), as well as between ambiguity and uncertainty (Kim, 2025; Cheng and Amiri, 2024), however, these papers construe uncertainty as a signal for poor performance and ambiguity as low quality inputs, which differs from our definition of ambiguity as an indispensable feature of language. Work in Question Answering has found that the best methods for detecting ambiguous inputs involve quantifying repetition within sampled model outputs (Cole et al., 2023), and using white-box metrics such as entropy (Yang et al., 2025).
UQ in MT has been used as a proxy for Quality Estimation (QE). For example, Fomicheva et al. (2020) use MT model uncertainty to estimate translation quality without references, while Glushkova et al. (2021) apply the same technique to the uncertainty of the QE models themselves. Other approaches use UQ to identify difficult instances and enhance training by applying curriculum learning (Zhou et al., 2020), semantic augmentation (Wei et al., 2020), balancing of multilingual training data (Wu et al., 2021) or test-time adaptation (Zhan et al., 2023). Wang et al. (2024b) examine zero-shot translation and distinguish between model uncertainty and data uncertainty, however their focus with regard to data uncertainty is on noisy, low-quality training data rather than inherent ambiguity. Cognitive science research has demonstrated that entropy-based uncertainty metrics are a suitable measure of ambiguity in human translations Bangalore et al. (2016), but this insight has yet to be applied in MT. To the best of our knowledge, no prior UQ-based approach in MT has explored ambiguity as a particular type of data uncertainty.
3 Method
We propose to quantify gender bias in Neural Machine Translation (NMT) models by characterising how gender is assigned to nouns across the predictive distribution. To do so, we base our methods on recently proposed UQ metrics which are founded on the classic Shannon entropy but take into account similarities between random Monte Carlo samples from the model. We first provide a brief overview of these UQ methods.
Let $\mathcal{Y}$ be a random variable whose value is drawn from the predictive distribution of an NMT model $p(y|x)$. Then entropy is defined as:
$\mathcal{H}(\mathcal{Y})=\underset{y\sim\mathcal{Y}}{\mathbb{E}}\left[I(y)\right],$
where $I$ is the surprisal of $y$. In the classic Shannon entropy, $I=-\log p(y)$, but the UQ methods we consider vary in their definition of surprisal.
Semantic Entropy (se; Farquhar et al., 2024) identifies semantic equivalences between elements and clusters them together according to a textual entailment model, mapping each $y$ to a cluster $c$. In our implementation we use a multilingual mDeberta model (He et al., 2021) finetuned on the Natural Language Inference (NLI) task by Laurer et al. (2022). Then, surprisal is the negative log probability of an element being in $c$:
$I_{\mathrm{SE}}(y)=-\log\underset{y^{\prime}\sim\mathcal{Y}}{\mathbb{E}}1\left[y^{\prime}\in{c}\right].$
Similarity-sensitive Shannon Entropy (s3e; Ricotta and Szeidl, 2006; Cheng and Vlachos, 2024) sets the surprisal of $y$ to the negative log of its expected similarity with all other outputs:
$I_{\mathrm{S3E}}(y)=-\log\mathbb{E}_{y^{\prime}\sim\mathcal{Y}}\left[\mathcal{S}(y,y^{\prime})\right],$
where $\mathcal{S}$ is a similarity function satisfying $\mathcal{S}(y,y^{\prime})\in[0,1]$ and $\mathcal{S}(y,y^{\prime})=1$ if $y=y^{\prime}$. Following Cheng and Vlachos (2024), we use cosine similarity between sentence embeddings of $y$ and $y^{\prime}$ generated by a multilingual E5 text embedding model (Wang et al., 2024a).
We also define Gender Entropy (ge), which is calculated like se but clusters elements based on the gender class of the translated focus noun. To determine the gender class, we use Spacy [^2] https://spacy.io/ and pymorphy2 (Korobov, 2015) morphological parsers.
se, s3e, and ge must be approximated with random sampling from $p(\cdot|x)$, which we perform with $\epsilon$-sampling (Hewitt et al., 2022), drawing 128 samples per source sentence. Further details about these UQ methods are given in Appendix A.
Our gender bias metrics are based on surprisal and entropy given by these UQ methods. The first desideratum is that for source sentences with unambiguous gender, an unbiased model should have lower surprisal of a translation with correct gender inflection compared to an incorrect inflection. Therefore, unbiased models should minimise relative surprisal, defined as:
$\Delta I=\frac{I(y_{\text{correct}})-I({y_{\text{incorrect}}})}{\frac{1}{2}(I({y_{\text{correct}})}+I({y_{\text{incorrect}}})}.$
The second desideratum is that the entropy of unbiased models should not be affected by the presence of bias cues. Thus, we define normalised entropy, which compares the $\mathcal{H}$ of a source sentence $x$ to the average entropy across its contrast set $\mathcal{G}_{x}$. $\mathcal{G}_{x}$ is a group of minimally different sentences that are identical to $x$ apart from the pronoun (e.g., ‘she’, ‘he’, ‘they’), including $x$ itself. The three sentences in Table 1 comprise a $\mathcal{G}_{x}$. Formally:
[…]
This formulation isolates variation in $\mathcal{H}(\mathcal{Y}_{x})$ attributable specifically to gender by holding all other lexical, syntactic, and semantic content constant across the contrast set.
The third desideratum is that models should show higher uncertainty for an input which is ambiguous with regard to gender, as compared to an input which is unambiguous, disregarding all biases in the input. It should therefore minimise relative entropy, defined as:
$\Delta\mathcal{H}=\frac{\mathcal{H}(\mathcal{Y}_{\text{unambiguous}})-\mathcal{H}(\mathcal{Y}_{\text{ambiguous}})}{\frac{1}{2}\left(\mathcal{H}(\mathcal{Y}_{\text{unambiguous}})+\mathcal{H}\left(\mathcal{Y}_{\text{ambiguous}}\right)\right)}.$
4 Experimental setup
To test our proposed bias metrics, an MT dataset containing information about gender ambiguity and stereotypes in the source sentences is required. We use WinoMT (Stanovsky et al., 2019), which includes annotations on minimal pairs of 1,584 sentences with masculine, feminine or neutral pronouns referring back to stereotypical or anti-stereotypical gender roles. An example of three sentences from the dataset can be seen in Table 1. The gender of the focus noun ‘mechanic’ is unambiguous in the first two sentences based on the contextual information (Pronoun M & F), but remains ambiguous in the third on account of the neutral (N) pronoun. The Stereotype (column 3) that mechanics are more often men than women either contradicts (row 1) or aligns with (row 2) the disambiguating context (column 2).
Sentence Pronoun Stereotype Subject Recency IC
The mechanic called to inform someone that he had completed the repair. M M M N N
The mechanic called to inform someone that she had completed the repair. F M F N N
The mechanic called to inform someone that they had completed the repair. N M N N N
Table 1: WinoMT Stanovsky et al. (2019) examples with additional annotations of bias cues.
This dataset also contains other linguistic phenomena that were not explicitly annotated in its released version. Thus, we automatically annotate additional linguistic bias cues, namely subject, recency, implicit causality, and person names, using syntactic parses with Spacy2. We release the additional annotations to the public for reproducibility.
For the Subject bias, following the literature on human biases for coherence (Nieuwland and Van Berkum, 2006), we hypothesise that models may assume coreference between the subject of the main clause (often the focus noun) and the subject of the complement clause (often the pronoun). In the example in Table 1the mechanic’ is the subject, therefore the Subject bias primes an interpretation in which the gender of the subject aligns with that of the subsequent pronoun (M in row 1, F in row 2). Furthermore, person names have been shown to have a strong effect on pronoun resolution (Saunders and Olsen, 2023). To assess the impact of person names on gender translation, we augment the dataset with common feminine and masculine names matching the gender of the pronoun, selected for their cross-linguistic familiarity (see Appendix B). As an example, when translating into French, the second sentence from Table 1 would read “The mechanic Anne called to inform someone that she had completed the repair."
Recency bias elicits the attribution of the gender of the most recent noun phrase to the following pronoun (Gautam et al., 2024). However, in our example, the most recent noun phrase is ‘someone’, so the focus noun is unaffected. The implicit causality (IC) bias induces an expectation in humans and LMs that when an IC verb is followed by a causal connective, the following pronoun will corefer with either the Subject or the Object of the IC verb, depending on the causal inference (Kementchedjhieva et al., 2021). For instance, a clause with the verb ‘call’, when followed by an explanation starting with ‘because’, is expected to be followed by a mention of the caller rather than the callee. We use an IC verb corpus (Garnham et al., 2021) for the annotation and find that about 10% of the dataset contains IC structures. The sentences in Table 1 do not include a causal complement, hence the IC bias does not apply.
Target Languages
We selected target languages that express gender through morphological markers on nouns and adjectives, and sometimes verbs, namely Spanish, French, Ukrainian, and Russian. The target languages vary in their representation within NLP research. A relevant idiosyncrasy of the Russian language is that for some nouns describing professions, even if a feminine version exists, it may be considered derogatory in use (Komova, 2024). For example, ‘врач’ is the masculine term for ‘doctor’, and the alternative feminine ‘врачиха’ is considered rude, which leads to ‘врач’ being used even when the doctor is known to be a woman. This results in constructions where masculine nouns are paired with feminine verb forms, or masculine markers are used throughout the sentence. In order to account for this, we also include a classification of professions which adhere to such (lack of) gender marking into WinoMT, using data from Komova (2024).
Human Translations
WinoMT does not contain target translations, thus the accuracy of MT models cannot be directly evaluated. To overcome this, we hired professional translators to translate a set of 100 WinoMT sentences into French, Spanish, Ukrainian and Russian. Each sentence is translated twice, with the focus noun in feminine and masculine variants respectively. They also annotate the translations as Correct or Incorrect with regard to the gender translation in the given context. For instance, when translating the English sentence “The farmer bought a book from the writer and paid her" into French, where ‘writer’ is the focus noun, the feminine ‘l’auteure’ should be marked as Correct, while ‘l’auteur’ would be Incorrect. In ambiguous cases, i.e. if the pronoun in the above sentence was ‘they’, both gender translations would be Correct. Appendix C provides the translation guidelines and details, and Appendix D discusses the quality of human annotations. We release the translations and correctness annotations to the public to enable further research.
Models
We experiment with two commonly used translation models, namely Opus-MT (Tiedemann and Thottingal, 2020) and m2m100 (Fan et al., 2021). Opus-MT models are NMT models trained on freely available parallel corpora. m2m100 is a many-to-many multilingual translation model, which directly translates between any pair of 100 languages. To examine how effective debiasing is regarding the three desiderata stated in Section 3, we apply the hard-debiasing method from Iluz et al. (2024) on the Opus-MT models, which have been shown to lower bias scores on the WinoMT dataset (Stanovsky et al., 2019) while maintaining translation quality. The hard-debiasing method neutralises the biased words in the representation space, so that neutral words are not associated with a specific gender (Bolukbasi et al., 2016). We adopt the most effective debiasing approach from Iluz et al. (2024), which applies debiasing to one-token profession words on the encoder side. See Appendix E for the performance of all models.
5 Research questions
Does Semantic Uncertainty Capture Gender Bias?
To validate the application of UQ metrics for bias evaluation, we compare their scores with the established gender accuracy metric. Gender accuracy uses the morphological parsers described in Section 3 to determine the focus noun gender in translations. As it relies on gold-standard references, it is applicable only to unambiguous items and unsuitable for cases with multiple valid gender realisations. We therefore limit this experiment to unambiguous items. We rank all models according to their $\Delta I$ scores and compare this ranking to that based on gender accuracy using Kendall’s $\tau$ and Pearson’s $r$. In order to establish whether a sampling-based metric is necessary, we also test a simple $\Delta LogProb$ value, which compares the Log Probabilities assigned to correct and incorrect instances, as an alternative to $\Delta I$.
What Biases do Models Exhibit via Uncertainty?
To evaluate model bias, we assess how bias cues influence the diversity of gender markers in translations. Specifically, we perform an analysis of variance (anova) to examine the effect of bias cues (independent variables) on normalised entropy measures $-\mathcal{H}$ (s3e), $-\mathcal{H}$ (se), and $-\mathcal{H}$ (ge) (dependent variables), using T-tests for significance. This analysis incorporates linguistic biases not previously explored in WinoMT.
What does Semantic Entropy Reveal about Bias with Ambiguity?
To evaluate model bias in ambiguous settings, which has not yet been explored, we compare models using their $\Delta\mathcal{H}$ scores. To isolate the uncertainty caused by ambiguity from that resulting from poorer model performance, we analyse the relationship between the $\Delta\mathcal{H}$ scores and the translation quality measured by the comet metric (Rei et al., 2022). Since the WinoMT dataset does not contain gold target translations, we use the 100 professionally annotated items described in Section 4, and the WMT test sets which contain the target languages (Callison-Burch et al., 2012; Bojar et al., 2013, 2014; Koehn et al., 2023; Haddow et al., 2024) for translation quality evaluation.
6 Results
This section presents the results of the bias evaluation using semantic uncertainty metrics.
Lang. Model Gender Acc ΔLog $\Delta\text{Log prob}$ $\Delta I$ (s3e) comet ES Opus-MT 67.95 0.00 -0.10 84.90 deb-Opus-MT 68.13 0.00 -0.13 84.86 m2m100 70.77 0.00 -0.13 72.05 FR Opus-MT 64.27 0.01 -0.04 83.56 deb-Opus-MT 64.79 0.01 -0.08 83.55 m2m100 61.66 0.01 -0.07 73.06 UK Opus-MT 45.34 0.00 -0.03 70.79 deb-Opus-MT 46.12 0.00 -0.03 70.79 m2m100 47.76 0.00 -0.02 52.85 RU Opus-MT 48.57 0.00 0.00 79.37 deb-Opus-MT 48.42 0.00 -0.03 79.36 m2m100 48.49 0.00 -0.03 58.62
Table 2: Gender Accuracy, $\Delta$ Log Probability, and $\Delta I$ (s3e) on Unambiguous instances, comet scores on WMT test sets (see Appendix E for details).
Model Rankings According to Semantic Surprisal and Gender Accuracy Correlate.
The results of the first experiment are presented in Table 2. We find that while ΔLog $\Delta\text{Log prob}$ does not correlate with the gender accuracy ranking, $\Delta I$ (s3e) shows a statistically significant negative correlation with gender accuracy (Kendall’s $\tau=-0.58$, Spearman’s $\rho=-0.78$; see Appendix K). We attribute the effectiveness of this metric in distinguishing Correct from Incorrect gender translations to its ability to capture gender information beyond noun morphology, including verb inflections and agreement, through its embedding representations. The flexibility of s3e enables it to encode nuances that ge does not. For example, when translating a sentence with a feminine pronoun into Russian, Opus-MT generates sentences with a masculine noun and a verb that is either feminine or masculine (e.g., “Перевозчик поблагодарила (fem) / поблагодарил (masc)" “The courier thanked"). This is reflected in a higher $\mathcal{H}$ score by s3e (0.65) than ge (0.00), as the variation in verb inflections is only captured by s3e.
The strong negative correlation between $\Delta I$ (s3e) and gender accuracy thus validates the core component of our proposed metric for evaluating bias in machine translation.
In addition, the ranking of models by their overall performance, as indicated by comet scores, is partially aligned with the rankings based on Gender Accuracy and $\Delta I$ for unambiguous instances. This suggests that in these instances, better performing models tend to be less biased (all rankings are listed in Appendix J).
Lang. Model Names Recency Implicit Causality Stereotype Subject Pronoun Default M Ambiguity
F M S F S M O F O M S F S M O F O M F M S F S M O F O M S O
es Opus-MT 0.41 0.41 -0.05 0.25 -0.19 0.24 -0.33 0.06 0.10 0.13 0.17 0.38 -0.21 0.24 -0.31 0.29 -0.13 N/A N/A -0.18
deb-Opus-MT -0.05 0.30 -0.11 0.27 -0.20 0.16 -0.38 0.05 0.14 0.08 0.05 0.49 -0.14 0.39 -0.24 0.14 -0.21 N/A N/A -0.10
m2m100 0.14 0.33 -0.11 0.29 -0.42 0.28 -0.25 0.18 0.27 0.12 0.07 0.51 -0.04 0.52 -0.08 0.04 -0.35 N/A N/A -0.11
fr Opus-MT 0.54 0.42 0.16 0.05 -0.34 0.06 -0.24 0.43 0.45 0.26 0.21 0.16 -0.14 0.16 -0.17 0.20 -0.04 N/A N/A -0.29
deb-Opus-MT 0.19 0.22 0.02 0.05 -0.25 0.17 -0.11 0.23 0.32 0.23 0.13 0.24 -0.03 0.31 -0.08 0.01 -0.14 N/A N/A -0.12
m2m100 -0.12 0.11 -0.23 0.50 0.09 0.49 0.03 -0.03 0.11 0.21 0.08 0.70 0.31 0.47 0.05 -0.08 -0.40 N/A N/A 0.06
uk Opus-MT -0.27 0.00 -0.11 0.26 -0.02 0.19 0.13 -0.11 0.17 0.03 -0.14 0.27 0.21 0.22 0.13 -0.11 -0.23 N/A N/A 0.06
deb-Opus-MT -0.43 -0.06 -0.12 0.41 0.00 0.18 0.07 -0.24 0.01 0.01 -0.17 0.23 0.17 0.16 0.10 -0.11 -0.17 N/A N/A 0.09
m2m100 0.08 -0.04 -0.18 0.27 0.02 0.33 0.18 0.19 0.37 -0.03 -0.17 0.53 0.34 0.50 0.26 -0.30 -0.42 N/A N/A 0.11
ru Opus-MT 0.01 -0.28 -0.41 0.00 -0.12 0.08 -0.10 -0.19 -0.20 -0.41 -0.39 0.07 -0.03 0.14 0.03 -0.10 -0.24 -0.40 0.04 0.35
deb-Opus-MT 0.23 -0.10 -0.17 0.05 0.03 0.05 -0.03 -0.10 -0.04 -0.11 -0.14 0.09 0.04 -0.13 0.03 -0.08 -0.13 -0.26 0.00 0.13
m2m100 -0.74 -0.16 -0.21 0.10 -0.04 -0.12 -0.25 0.12 0.12 -0.20 -0.18 0.06 -0.02 0.08 0.01 -0.07 -0.11 -0.20 -0.22 0.18
Table 3: anova results: single effects of bias cues (Feminine, Masculine, Subject and Object) on $-\mathcal{H}$ (s3e). Values correspond to effect coefficients (deviations from a reference group). Boldface indicates statistical significance (p < 0.05). The sign of the values indicates whether the presence of the variable increases (positive) or decreases (negative) the mean $\mathcal{H}$ of the group containing the given variable value. Reference group is N for all columns except: ‘no name’ for Names, ‘no default’ for Default M, ‘unambiguous’ for Ambiguity.
Semantic Entropy Scores Vary With Regard to Bias Cues.
The results of the second experiment presented in Table 3 show that most bias cues in the data have a significant effect on the variance of $-\mathcal{H}$ (s3e), indicating that the tested models exhibit various social and linguistic biases.333 $-\mathcal{H}$ (s3e) shows the strongest sensitivity to bias cues compared to $-\mathcal{H}$ (se) and $-\mathcal{H}$ (ge). We also experiment with unnormalised $\mathcal{H}$ scores, the results of which are less comparable across metrics, bias types, and models. The full results are presented in Appendix F. All trends observed for $-\mathcal{H}$ (s3e) are also present in $-\mathcal{H}$ (se) and $-\mathcal{H}$ (ge), as well as for unnormalised $\mathcal{H}$. The results corroborate previous findings. The high absolute coefficient values in the Names column indicate that person names have an effect on gender translation even when a disambiguating pronoun is present. This is in line with Saunders and Olsen (2023), who have shown that both pronouns and names induce gender bias and are often not sufficient for full disambiguation. Secondly, the fact that some Russian nouns have a default masculine grammatical gender regardless of the context (Komova, 2024) is reflected in significant decrease in gender diversity for sentences containing such nouns (negative coefficients in the Default M columns indicating lower $-\mathcal{H}$). Thirdly, we observe that masculine biases generally reduce $-\mathcal{H}$ (negative coefficients in the M columns), while feminine biases tend to increase it (positive coefficients in the F columns). This suggests a general default toward masculine translations in models, with outputs becoming more similar under masculine biases and more varied under feminine ones. This finding aligns with Kuzucu et al. (2025), who show that model uncertainty is typically higher for minority groups.
Translation Accuracy Affects the Bias–Entropy Relationship Differently Across Levels of Analysis.
Having validated the s3e metric in the first experiment, we investigate the results of $\Delta\mathcal{H}$ (s3e) as a bias metric. Table 4 illustrates that some models (Opus-MT-UK, deb-Opus-MT-UK, m2m100-UK, Opus-MT-RU, m2m100-RU) exhibit the desired negative $\Delta\mathcal{H}$. Surprisingly, this result suggests that when it comes to ambiguous instances, contrary to the unambiguous cases, the models which perform better on translation accuracy (namely models for Spanish and French) are not generally less gender biased (model rankings according to all metrics are available in Appendix J). This finding mirrors the results in the Ambiguity column in Table 3, where $-\mathcal{H}$ increases for Ukrainian and Russian (positive coefficients), but not for Spanish and French. Higher $-\mathcal{H}$ for Ambiguous items (or negative $\Delta\mathcal{H}$) is expected for an unbiased model.
Lang. Model Unamb Amb $\Delta\mathcal{H}$
ES Opus-MT 1.23 1.12 0.09
deb-Opus-MT 0.97 0.89 0.08
m2m100 1.79 1.45 0.19
FR Opus-MT 1.79 1.43 0.20
deb-Opus-MT 1.21 1.08 0.11
m2m100 3.22 2.78 0.14
UK Opus-MT 1.96 2.16 -0.10
deb-Opus-MT 1.98 2.15 -0.09
m2m100 2.05 2.28 -0.11
RU Opus-MT 1.56 1.68 -0.08
deb-Opus-MT 1.05 0.97 0.08
m2m100 1.83 2.29 -0.25
Table 4: Unambiguous and Ambiguous $\mathcal{H}$ (s3e)
In contrast, we observe the expected effect of debiasing on the Spanish and French models (lower $\Delta\mathcal{H}$ for deb-Opus-MT in Table 4), suggesting that models which perform better overall are more susceptible to debiasing. The impact of debiasing is also reflected in Table 3, as the effects are mostly smaller (lower absolute values of coefficients) in debiased models compared to their non-debiased counterparts across languages, confirming that debiasing is at least partially effective.
Figure 2: Violin plots of binned comet scores and $\mathcal{H}$ (s3e) on ambiguous and unambiguous inputs. Low, medium and high comet scores from left to right, evaluated with human translations, multi-reference for ambiguous items.
In Figure 2, results are grouped by comet score bins for a more fine-grained analysis at the instance level. For the models with negative $\Delta\mathcal{H}$ (s3e) scores (Ukrainian and Russian), $\Delta\mathcal{H}$ is typically most pronounced for the highest-accuracy translations (e.g. ambiguous scores for m2m100-RU in Bin 3 are substantially higher than B1). Although debiasing does not reduce the overall $\Delta\mathcal{H}$ score for Ukrainian (see Table 4), it results in the largest improvement in the highest-quality translations: in bin B3, ambiguous $\mathcal{H}$ scores for deb-Opus-MT-UK are notably higher than those of the original model. This improvement is further supported by a substantial 8.41% drop in masculine focus noun inflections for Ukrainian, compared to 0.88–2.49% for other languages. We hypothesise that this is due to the limited training data in Ukrainian, which may lead to a less stable model that performs worse overall but is more responsive to debiasing in high-quality outputs. The relationship between translation accuracy and bias under ambiguity appears to differ depending on the level of analysis: between models, higher accuracy does not imply lower bias on ambiguous instances, whereas within models, higher-accuracy instances tend to show lower bias.
Qualitative Analysis
A qualitative analysis of the example in Table 1, presented with corresponding $\mathcal{H}$ values in Table 5, corroborates the quantitative findings in Figure 2. When translating the sentence in row 1, across all target languages, Opus-MT models consistently produce only the masculine variants of the focus noun (‘El mecánico’, ‘Le mécanicien’, ‘Механик’ and ‘Механiк’). In the anti-stereotypical case (row 2), all languages except Russian include both masculine and feminine forms (‘La mecánica’, ‘La mécanicienne’ and ‘Механiка’), indicating that these models are sensitive to the masculine stereotype even when the referent in the context is clearly feminine. For Russian, the models fail to generate any feminine constructions, even when the context is unambiguously feminine. This difference is evident in the $\mathcal{H}$ (s3e) scores, which are higher in row 2 than row 1 for the first three languages. Moreover, in the ambiguous case (row 3), all Opus-MT models produce only masculine nouns, regardless of language. Consequently, $\mathcal{H}$ (s3e) scores are generally higher for the unambiguous cases (mean of rows 1 and 2) than for the ambiguous case (row 3), except for Russian, where $\mathcal{H}$ remains low across all conditions. These observations are consistent with expectations for biased models: they default to stereotypical gender realisations when the pronoun is ambiguous and sometimes even when the context clearly suggests an anti-stereotypical interpretation.
Sentence ES $\rightarrow$ deb FR $\rightarrow$ deb UK $\rightarrow$ deb RU $\rightarrow$ deb The mechanic called to inform someone that he had completed the repair. 0.75 0.82 0.00 0.00 2.41 1.57 0.33 0.31 The mechanic called to inform someone that she had completed the repair. 1.64 1.85 0.53 0.56 2.75 2.05 0.00 0.00 The mechanic called to inform someone that they had completed the repair. 0.74 0.87 0.00 0.00 2.38 2.03 0.37 0.37
Table 5: WinoMT examples with $\mathcal{H}$ (s3e) values, for Opus-MT (left) and deb-Opus-MT (right) models.
The qualitative analysis also reveals interesting effects of debiasing. In the anti-stereotypical case in row 2, debiasing increases the number of feminine constructions generated in Spanish (from 43/128 to 55/128, corresponding to a slight increase in $\mathcal{H}$, as feminine forms remain a minority) and Ukrainian (from 72/128 to 128/128, reflected in a decrease in $\mathcal{H}$). No notable changes are observed for French or Russian, consistent with stable $\mathcal{H}$ (s3e) scores. When it comes to the ambiguous pronoun (row 3), the debiased models continue to generate only masculine variants of ‘mechanic’ in Spanish, French and Russian, with $\mathcal{H}$ remaining largely unchanged. In contrast, all debiased model outputs in Ukrainian include a feminine translation of the noun, corresponding to a decrease in $\mathcal{H}$ from Opus-MT to debiased Opus-MT in row 3. This pattern illustrates that when debiasing leads to overgeneration of feminine morphology in ambiguous contexts, our proposed metric flags this as increased bias (positive $\Delta\mathcal{H}$), indicating that such changes are not deemed as improvements.
7 Conclusion
In this work, we apply distribution-level UQ to evaluate bias in MT models. This method complements gender accuracy, particularly where gender accuracy is inapplicable. Specifically, it captures the more subtle manifestations of gender bias that arise when models show a preference for one gender in ambiguous contexts. Our overall contribution is the novel use of UQ as a bias metric in MT, which 1) does not rely on gender references, 2) is general and captures multiple types of bias, 3) is validated by the established metric of gender accuracy, and 4) provides new insights into biased behavior in ambiguous contexts, a setting not previously studied. Future work will extend the proposed bias evaluation method to tasks beyond translation.
Limitations
This study is limited to Romance and Slavic languages, not including many other language families which mark gender and express stereotypes in diverse ways. While we tried to account for language differences by including different names for different target languages, accounting for specific masculine-only nouns in Russian, debiasing with language-specific vocabularies, etc., some linguistic idiosyncracies are still not accounted for, such as the fact that profession stereotypes are defined in English and may apply differently in different regions. Finally, our work is limited to two grammatical genders, and treats ‘they’ as a neutral pronoun that may refer to any gender, however we do not study the interpretation of the pronoun as referring to non-binary people specifically. Further directions include applying UQ to ambiguity detection, which could enable more gender-inclusive translations through morphological doubling, where both masculine and feminine morphemes are included for gender neutrality. Future work should address these directions.
Ethics Statement
The models used in this study, like all ML models, can be biased and make mistakes, including in gender attribution. Our contribution aims to specifically tackle masculine and feminine gender stereotypes via more stringent evaluation metrics, in order to avoid the perpetuation of gender bias.
Acknowledgments
Ieva Raminta Staliūnaitė is supported by Huawei. Andreas Vlachos is supported by the ERC grant AVeriTeC (GA 865958). Julius Cheng is supported by a research grant from Imminent. The translations for the datasets were funded by the same Imminent research grant and we would like to thank Andrii Nikolaiev, Mykhailo Khokhlovych, Sarah Mokadem and Daniel Carter for their work on the translations.
Appendix A Further UQ discussion
Farquhar et al. (2024) does not define a per-element surprisal; the original definition computes Shannon entropy over clusters. $\mathcal{Y}$ are mapped to clusters $\mathcal{C}$, and SE is:
$\mathcal{H}_{\text{SE}}(\mathcal{C})=-\underset{c\sim\mathcal{C}}{\mathbb{E}}\log p(c|x).$
Surprisal is thus defined for a cluster instead of an element, but it is easy to show that our per-element surprisal obtains equivalent entropy as the original definition.
Cheng and Vlachos (2024) introduce a hyperparameter $\alpha$ which is applied as an exponent to the similarity function. This is used to rescale $\mathcal{S}$ for more favorable performance on benchmarks. We tune $\alpha$ for the highest correlation between s3e and the entropy of the gender labels assigned to the nouns in question by the morphological parser. This way we aim for $\mathcal{H}$ (s3e) to capture gender variation, with higher values of $\mathcal{H}$ (s3e) indicating more diversity in the gender morphemes.
We also experiment with alternative similarity metrics to cosine similarity, such as Euclidean Distance, Chebyshev Distance, Manhattan Distance, Minkowski Distance, etc. but find no notable differences and use the simplest option of Cosine Similarity in the paper.
Appendix B Annotation of Names
Table 6 presents the names used for expanding the WinoMT dataset to include common names for masculine and feminine genders in French, Spanish, Ukrainian and Russian.
Spanish French Ukrainian Russian
F Carla Anne Anna Anna
M Gabriel Victor Ivan Ivan
Table 6: The names used for extending the WinoMT dataset for different target languages.
Appendix C Translation Guidelines and Translator Recruitment Details
Professional translators were hired and paid £0.2 per word as requested. The translation guidelines are as follows:
Please translate the following sentences into [French/Spanish/Ukrainian/Russian]. Please provide two translations for each sentence, with different gender inflections for the highlighted noun —one masculine and one feminine. Please keep the pronoun true to the original to the extent that it is possible. Please use the translations that appear natural to you as a native speaker rather than using forms that are prescribed [^4] This instruction was added due to inquiries by translators about the use of formal or dominant dialects as well as language policy changes influencing how they might translate some words.. Please note whether both translations are acceptable in terms of the gender of the noun matching the contextual information, and if they are not, which one is the correct one:
  • The developer argued with the designer because she did not like the design.
  • The undergraduate presented the administrator with a petition that they had organized.
Appendix D Human Annotation Quality
The annotations of Correct and Incorrect sentences by professional translators are compared to the gold standard gender annotations in the dataset, and across all four translators range between 93.17 and 93.27 in Cohen’s $\kappa$ scores, confirming that apart from some linguistic idiosyncrasies of each language (e.g. ‘victim’ in Spanish is always feminine and so regardless of the contextualising pronoun will take the same form), the annotators agree on which sentences should be correctly translated in which gender.
Appendix E Overall Model Performance
Table 7 presents the performance of the models used in this study in terms of the comet metric [^5] https://huggingface.co/Unbabel/wmt22-comet-da (Rei et al., 2022) on WMT datasets which contain the target languages (Callison-Burch et al., 2012; Bojar et al., 2013, 2014; Koehn et al., 2023; Haddow et al., 2024). The models are run on a single NVIDIA TU102 GPU.
Model Opus-MT deb-Opus-MT m2m100
Dataset es fr uk ru es fr uk ru es fr uk ru
newstest2012 84.52 82.21 84.47 82.22 71.25 71.97
newstest2013 85.28 83.45 85.24 83.44 72.84 72.57
newstest2014 85.01 87.44 85.00 87.44 74.63 72.29
wmttest2023 74.58 79.02 74.58 79.02 56.49 55.93
wmttest2024 66.99 71.64 67.00 71.63 49.20 47.63
mean 84.90 83.56 70.79 79.37 84.86 83.55 70.79 79.36 72.05 73.06 52.85 58.62
Table 7: comet scores on WMT test sets for the models used.
Appendix F anova Results
Table 8 presents the anova results for s3e, se and ge metrics. Table 9 presents the anova results without normalising the $\mathcal{H}$ values.
Lang. Model Names Recency Implicit Causality Stereotype Subject Context Default M Ambiguity
F M S F S M O F O M S F S M O F O M F M S F S M O F O M S O
s3e
es Opus-MT 0.41 0.41 -0.05 0.25 -0.19 0.24 -0.33 0.06 0.10 0.13 0.17 0.38 -0.21 0.24 -0.31 0.29 -0.13 N/A N/A -0.18
deb-Opus-MT -0.05 0.30 -0.11 0.27 -0.20 0.16 -0.38 0.05 0.14 0.08 0.05 0.49 -0.14 0.39 -0.24 0.14 -0.21 N/A N/A -0.10
m2m100 0.14 0.33 -0.11 0.29 -0.42 0.28 -0.25 0.18 0.27 0.12 0.07 0.51 -0.04 0.52 -0.08 0.04 -0.35 N/A N/A -0.11
fr Opus-MT 0.54 0.42 0.16 0.05 -0.34 0.06 -0.24 0.43 0.45 0.26 0.21 0.16 -0.14 0.16 -0.17 0.20 -0.04 N/A N/A -0.29
deb-Opus-MT 0.19 0.22 0.02 0.05 -0.25 0.17 -0.11 0.23 0.32 0.23 0.13 0.24 -0.03 0.31 -0.08 0.01 -0.14 N/A N/A -0.12
m2m100 -0.12 0.11 -0.23 0.50 0.09 0.49 0.03 -0.03 0.11 0.21 0.08 0.70 0.31 0.47 0.05 -0.08 -0.40 N/A N/A 0.06
uk Opus-MT -0.27 0.00 -0.11 0.26 -0.02 0.19 0.13 -0.11 0.17 0.03 -0.14 0.27 0.21 0.22 0.13 -0.11 -0.23 N/A N/A 0.06
deb-Opus-MT -0.43 -0.06 -0.12 0.41 0.00 0.18 0.07 -0.24 0.01 0.01 -0.17 0.23 0.17 0.16 0.10 -0.11 -0.17 N/A N/A 0.09
m2m100 0.08 -0.04 -0.18 0.27 0.02 0.33 0.18 0.19 0.37 -0.03 -0.17 0.53 0.34 0.50 0.26 -0.30 -0.42 N/A N/A 0.11
ru Opus-MT 0.01 -0.28 -0.41 0.00 -0.12 0.08 -0.10 -0.19 -0.20 -0.41 -0.39 0.07 -0.03 0.14 0.03 -0.10 -0.24 -0.40 0.04 0.35
deb-Opus-MT 0.23 -0.10 -0.17 0.05 0.03 0.05 -0.03 -0.10 -0.04 -0.11 -0.14 0.09 0.04 -0.13 0.03 -0.08 -0.13 -0.26 0.00 0.13
m2m100 -0.74 -0.16 -0.21 0.10 -0.04 -0.12 -0.25 0.12 0.12 -0.20 -0.18 0.06 -0.02 0.08 0.01 -0.07 -0.11 -0.20 -0.22 0.18
se
es Opus-MT -1.64 0.19 -0.08 0.13 -0.19 0.09 -0.29 -0.01 0.02 0.06 0.07 0.28 -0.12 0.17 -0.22 0.15 -0.08 N/A N/A -0.05
deb-Opus-MT -0.06 0.27 0.13 -0.01 -0.07 0.00 -0.19 0.06 0.10 0.16 0.12 -0.01 -0.18 -0.01 -0.19 0.20 0.08 N/A N/A -0.20
m2m100 -1.71 0.23 -0.04 0.20 -0.29 0.17 -0.16 0.10 0.18 0.15 0.08 0.28 -0.06 0.25 -0.09 0.07 -0.16 N/A N/A -0.10
fr Opus-MT -1.59 0.19 0.00 0.09 -0.25 0.05 -0.17 0.23 0.25 0.15 0.11 0.11 -0.12 0.10 -0.15 0.14 -0.04 N/A N/A -0.09
deb-Opus-MT -0.01 0.24 0.08 0.11 -0.17 0.04 -0.10 0.23 0.20 0.19 0.18 -0.06 -0.22 0.01 -0.15 0.18 0.03 N/A N/A -0.16
m2m100 -0.15 0.29 0.16 0.03 -0.05 0.02 -0.14 0.27 0.24 0.27 0.30 0.01 -0.14 0.01 -0.15 0.19 0.06 N/A N/A -0.22
uk Opus-MT -1.54 0.00 -0.12 0.21 -0.05 0.10 0.02 -0.09 0.05 0.01 -0.08 0.13 0.04 0.10 0.02 0.00 -0.14 N/A N/A 0.06
deb-Opus-MT 0.09 0.12 -0.06 0.13 -0.07 0.08 0.01 -0.02 0.08 0.15 0.09 0.06 -0.11 -0.07 -0.10 0.13 0.06 N/A N/A 0.09
m2m100 -1.72 -0.04 -0.14 0.20 0.03 0.15 0.08 0.05 0.16 -0.02 -0.11 0.22 0.10 0.18 0.09 -0.09 -0.19 N/A N/A 0.09
ru Opus-MT -1.50 -0.17 -0.25 0.21 0.07 0.00 -0.02 -0.13 -0.18 -0.25 -0.28 0.03 -0.03 0.10 0.00 -0.06 -0.14 -0.32 -0.02 0.21
deb-Opus-MT -0.06 0.04 0.01 0.11 0.05 0.00 -0.06 -0.22 -0.13 0.07 -0.01 -0.11 -0.15 -0.05 -0.11 0.09 0.07 -0.24 -0.12 0.03
m2m100 -1.57 -0.09 -0.14 0.11 0.10 0.04 -0.18 -0.01 0.00 -0.10 -0.10 0.00 -0.06 0.03 -0.01 -0.01 -0.07 -0.12 -0.18 0.11
ge
es Opus-MT 0.02 0.30 0.04 -0.17 0.16 -0.10 -0.14 0.11 0.15 0.17 0.12 0.19 -0.12 0.19 -0.20 0.15 -0.06 N/A N/A 0.17
deb-Opus-MT -0.04 0.27 0.06 0.16 -0.10 0.15 -0.12 0.09 0.17 0.17 0.09 0.21 -0.09 0.24 -0.16 0.09 -0.06 N/A N/A -0.16
m2m100 -0.10 0.16 0.00 0.08 -0.10 0.09 -0.08 0.05 0.07 0.10 0.09 0.09 -0.08 0.08 -0.16 0.12 -0.01 N/A N/A -0.08
fr Opus-MT 0.01 0.20 0.05 0.12 -0.06 0.07 -0.11 0.09 0.13 0.12 0.08 0.09 -0.10 0.06 -0.19 0.15 0.03 N/A N/A -0.13
deb-Opus-MT -0.05 0.18 0.04 -0.11 0.08 0.08 -0.11 0.05 0.12 0.12 0.06 0.11 -0.09 0.10 -0.17 0.11 0.01 N/A N/A -0.11
m2m100 -0.02 0.19 0.02 0.12 -0.10 0.09 -0.12 0.08 0.10 0.11 0.07 0.08 -0.11 0.04 -0.19 0.17 0.02 N/A N/A -0.11
uk Opus-MT -0.04 0.05 -0.04 0.07 -0.03 0.03 -0.07 0.06 0.06 0.02 0.02 0.02 -0.06 -0.11 -0.15 0.16 0.06 N/A N/A 0.01
deb-Opus-MT -0.03 0.05 0.02 0.07 -0.03 0.00 -0.06 0.06 0.07 0.03 -0.03 -0.02 -0.05 -0.16 -0.17 0.17 0.14 N/A N/A 0.03
m2m100 -0.04 0.06 -0.04 0.09 -0.05 0.05 -0.06 0.07 0.08 0.01 0.01 0.03 -0.07 -0.07 -0.18 0.16 0.06 N/A N/A 0.01
ru Opus-MT -0.01 0.00 -0.03 -0.02 -0.01 0.00 -0.04 0.03 0.03 -0.04 -0.06 -0.01 -0.04 -0.11 -0.13 0.12 0.08 -0.04 0.00 -0.02
deb-Opus-MT -0.02 -0.01 -0.02 -0.05 -0.05 -0.01 -0.02 0.01 0.02 -0.04 -0.05 -0.03 -0.04 -0.15 -0.15 0.13 0.12 -0.05 -0.01 -0.02
m2m100 0.02 0.07 0.02 -0.01 -0.06 -0.01 -0.07 0.09 0.07 0.02 0.04 -0.01 -0.06 -0.05 -0.11 0.10 0.06 -0.04 -0.03 0.04
Table 8: anova results: single effects of bias cues (Feminine, Masculine, Subject and Object) on $-\mathcal{H}$ (s3e), $-\mathcal{H}$ (se) and $-\mathcal{H}$ (ge). Values correspond to effect coefficients (deviations from a reference group). Boldface indicates statistical significance (p < 0.05). The sign of the values indicates whether the presence of the variable increases (positive) or decreases (negative) the mean $\mathcal{H}$ of the group containing the given variable value. Reference group is N for all columns except: ‘no name’ for Names, ‘no default’ for Default M, ‘unambiguous’ for Ambiguity.
Lang. Model Names Recency Implicit Causality Stereotype Subject Context Default M Ambiguity
F M S F S M O F O M S F S M O F O M F M S F S M O F O M S O
s3e
es Opus-MT 12.12 47.0 83.53 -2.45 20.67 -2.44 27.76 38.98 49.12 44.16 37.6 0.23 31.36 7.95 42.89 -22.44 14.64 N/A N/A -65.29
deb-Opus-MT -2.35 36.52 79.64 -58.1 -7.06 18.24 -8.44 31.09 33.94 42.22 36.85 32.48 -1.86 34.19 6.06 48.72 -29.24 N/A N/A 14.06
m2m100 0.19 0.54 0.02 -0.28 0.38 -0.27 0.37 -0.26 0.18 0.25 0.25 0.17 0.31 -0.25 0.31 -0.29 0.31 N/A N/A -0.17
fr Opus-MT 0.12 0.41 0.06 0.32 -0.15 0.23 -0.17 0.18 0.25 0.23 0.14 0.26 -0.16 0.32 -0.19 0.16 -0.14 N/A N/A -0.24
deb-Opus-MT -0.02 0.26 0.06 -0.16 0.23 -0.07 0.15 -0.1 0.12 0.21 0.18 0.08 0.21 -0.08 0.25 -0.12 0.05 N/A N/A -0.08
m2m100 -0.12 0.11 -0.23 0.06 0.5 0.09 0.49 0.03 -0.03 0.11 0.21 0.08 0.7 0.31 0.47 0.05 -0.08 N/A N/A -0.4
uk Opus-MT 0.12 0.22 0.03 0.16 -0.07 0.09 -0.11 0.15 0.18 0.14 0.1 0.09 -0.1 0.05 -0.1 0.16 -0.04 -0.09 -0.05 -0.12
deb-Opus-MT 0.83 -10.2 24.93 -7.38 -19.32 18.37 -14.82 27.31 5.26 9.67 8.87 5.29 -15.13 23.85 -10.15 44.14 -26.98 1.7 -2.21 -0.34
m2m100 0.14 0.3 0.06 -0.18 0.2 -0.16 0.12 -0.13 0.19 0.19 0.16 0.15 0.1 -0.15 0.05 -0.17 0.22 -0.02 -0.1 -0.04
ru Opus-MT 15.38 4.07 52.95 -27.98 17.38 -24.19 20.45 11.79 12.51 12.13 13.6 -24.94 22.11 -28.52 41.31 -18.56 23.26 -0.79 -1.99 -28.53
deb-Opus-MT 0.23 -0.1 -0.17 0.13 0.05 0.03 0.05 -0.03 -0.1 -0.04 -0.11 -0.14 0.09 0.04 0.13 0.03 -0.08 -0.13 -0.26 -0.0
m2m100 0.08 -0.07 -0.05 0.06 -0.02 -0.01 -0.02 -0.0 -0.06 -0.07 -0.05 -0.03 -0.01 0.0 -0.02 0.0 -0.01 -0.0 0.01 0.03
se
es Opus-MT 12.12 47.0 83.53 -2.45 20.67 -2.44 27.76 38.98 49.12 44.16 37.6 0.23 31.36 7.95 42.89 -22.44 14.64 N/A N/A -65.29
deb-Opus-MT -0.08 -0.06 -0.05 0.05 -0.01 0.0 -0.01 -0.0 -0.04 -0.05 -0.04 -0.03 -0.01 0.0 -0.01 0.01 -0.02 N/A N/A -0.01
m2m100 0.08 -0.07 -0.06 0.07 -0.01 -0.0 -0.02 -0.01 -0.05 -0.05 -0.04 -0.04 -0.01 0.0 -0.02 0.01 -0.02 N/A N/A -0.01
fr Opus-MT 16.01 47.22 80.76 -0.14 19.01 -8.69 26.55 31.41 35.76 33.4 31.98 -1.93 29.87 2.67 35.97 -15.81 17.83 N/A N/A -64.01
deb-Opus-MT -0.07 -0.05 -0.04 0.04 -0.0 0.02 -0.01 0.0 -0.03 -0.04 -0.03 -0.03 -0.01 0.01 -0.01 0.01 -0.02 N/A N/A -0.01
m2m100 0.14 0.38 0.04 -0.21 0.35 -0.15 0.23 -0.17 0.15 0.2 0.19 0.13 0.23 -0.16 0.24 -0.19 0.19 N/A N/A -0.12
uk Opus-MT -0.01 -0.01 -0.01 -0.03 -0.02 0.0 0.0 -0.03 -0.03 -0.01 -0.01 -0.0 0.0 -0.01 0.01 -0.0 -0.0 0.02 0.03 0.01
deb-Opus-MT 0.0 0.09 0.02 -0.05 0.01 -0.06 0.06 -0.04 0.12 0.16 0.08 0.04 0.04 -0.02 0.08 0.02 0.01 -0.06 -0.1 -0.03
m2m100 0.14 0.3 0.06 -0.18 0.2 -0.16 0.12 -0.13 0.19 0.19 0.16 0.15 0.1 -0.15 0.05 -0.17 0.22 -0.02 -0.1 -0.04
ru Opus-MT 15.38 4.07 52.95 -27.98 17.38 -24.19 20.45 11.79 12.51 12.13 13.6 -24.94 22.11 -28.52 41.31 -18.56 23.26 -0.79 -1.99 -28.53
deb-Opus-MT 1.34 -3.62 46.99 -21.71 -33.31 16.73 -27.33 19.72 9.4 7.79 6.6 9.66 -28.51 22.12 -24.71 50.74 -26.37 15.86 -0.28 -0.68
m2m100 20.01 18.76 50.38 -34.59 -11.85 5.91 -16.63 13.63 14.94 15.4 14.21 15.2 -17.46 14.68 -15.4 30.42 -10.35 16.48 -0.74 -3.23
ge
es Opus-MT -0.0 -0.02 -0.01 -0.01 0.01 -0.01 0.0 -0.02 -0.02 -0.01 -0.01 -0.01 0.01 -0.0 0.01 -0.01 -0.0 N/A N/A 0.02
deb-Opus-MT -2.35 36.52 79.64 -58.1 -7.06 18.24 -8.44 31.09 33.94 42.22 36.85 32.48 -1.86 34.19 6.06 48.72 -29.24 N/A N/A 14.06
m2m100 0.19 0.54 0.02 -0.28 0.38 -0.27 0.37 -0.26 0.18 0.25 0.25 0.17 0.31 -0.25 0.31 -0.29 0.31 N/A N/A -0.17
fr Opus-MT 0.06 -0.05 -0.04 0.01 0.02 -0.01 0.01 -0.04 -0.04 -0.03 -0.02 -0.0 0.01 -0.01 0.01 -0.02 -0.01 N/A N/A 0.04
deb-Opus-MT -0.05 0.18 0.04 -0.11 0.08 -0.11 0.08 -0.11 0.05 0.12 0.12 0.06 0.11 -0.09 0.1 -0.17 0.11 N/A N/A 0.01
m2m100 0.14 0.38 0.04 -0.21 0.35 -0.15 0.23 -0.17 0.15 0.2 0.19 0.13 0.23 -0.16 0.24 -0.19 0.19 N/A N/A -0.12
uk Opus-MT 0.12 0.22 0.03 0.16 -0.07 0.09 -0.11 0.15 0.18 0.14 0.1 0.09 -0.1 0.05 -0.1 0.16 -0.04 -0.09 -0.05 -0.12
deb-Opus-MT 0.0 0.09 0.02 -0.05 0.01 -0.06 0.06 -0.04 0.12 0.16 0.08 0.04 0.04 -0.02 0.08 0.02 0.01 -0.06 -0.1 -0.03
m2m100 0.14 0.3 0.06 -0.18 0.2 -0.16 0.12 -0.13 0.19 0.19 0.16 0.15 0.1 -0.15 0.05 -0.17 0.22 -0.02 -0.1 -0.04
ru Opus-MT 0.09 0.17 0.05 0.01 -0.11 0.05 -0.08 0.08 0.13 0.11 0.06 0.04 -0.08 -0.02 -0.06 0.14 -0.01 -0.08 -0.03 -0.11
deb-Opus-MT -0.02 -0.01 -0.02 0.02 -0.05 -0.05 -0.01 -0.02 0.01 0.02 -0.04 -0.05 -0.03 -0.04 -0.15 -0.15 0.13 0.12 -0.05 -0.01
m2m100 0.15 0.29 0.08 -0.18 0.13 -0.1 0.09 -0.12 0.16 0.21 0.15 0.11 0.07 -0.13 0.08 -0.05 0.15 -0.08 -0.1 -0.02
Table 9: anova results (no normalisation): single effects of bias cues (Feminine, Masculine, Subject and Object) on $\mathcal{H}$ (s3e), $\mathcal{H}$ (se) and $\mathcal{H}$ (ge). Values correspond to effect coefficients (deviations from a reference group). Boldface indicates statistical significance (p < 0.05). The sign of the values indicates whether the presence of the variable increases (positive) or decreases (negative) the mean $\mathcal{H}$ of the group containing the given variable value. Reference group is N for all columns except: ‘no name’ for Names, ‘no default’ for Default M, ‘unambiguous’ for Ambiguity.
Appendix G Gender Accuracy
Table 10 presents more fine-grained results than Table 2 with regard to gender accuracy, namely splitting the results by subset of the dataset. The results in the Ambiguous column are not meaningfully interpretable, as a single ground truth label of gender cannot capture the true desired behavior of the model, especially when the gold label for ambiguous cases is mostly ‘neutral’, and neutral is not commonly used as grammatical gender for animate objects in the languages used in this study. The case of Russian, where the performance increases on the Ambiguous subset actually reflects the model choosing the masculine forms, which are tagged as ‘neutral’ by the morphological parser due to the masculine form often being the default choice for both genders, as discussed in Section 4.
Lang. Model All Pro Anti Unamb. Amb.
es Opus-MT 55.20 67.95 52.10 67.95 33.96
deb-Opus-MT 55.69 68.13 52.95 68.13 34.39
m2m100 55.68 70.77 51.17 70.77 32.40
fr Opus-MT 52.05 64.27 46.55 64.27 37.25
deb-Opus-MT 52.98 64.79 48.10 64.79 37.75
m2m100 50.95 61.66 47.57 61.66 34.84
uk Opus-MT 38.65 45.34 34.20 45.34 33.75
deb-Opus-MT 38.95 46.12 34.15 46.12 33.74
m2m100 40.97 47.76 36.81 47.76 35.20
ru Opus-MT 39.50 48.57 33.27 48.57 33.24
deb-Opus-MT 39.50 48.42 33.38 48.42 33.33
m2m100 41.01 48.49 36.81 48.49 33.81
Table 10: Comparison of Gender Accuracy Overall, in Pro-/Anti-Stereotypical and Ambiguous Cases Across Models, on WinoMT.
Appendix H Quality Estimation with Human Translations
Table 11 presents the results of the models used in this study on the 100 human-annotated instances. For unambiguous cases we use a single reference, whereas for ambiguous cases, we calculate performance by taking the maximum comet score of both acceptable translations.
Lang. Model All Pro Anti Unamb. Amb.
es Opus-MT 81.35 85.80 83.37 84.55 75.14
deb-Opus-MT 81.31 85.62 83.43 84.49 75.14
m2m100 79.56 84.44 81.29 82.82 73.25
fr Opus-MT 77.63 82.23 81.16 81.66 70.47
deb-Opus-MT 77.69 82.41 80.88 81.60 70.73
m2m100 76.24 80.69 78.65 79.61 70.25
uk Opus-MT 80.56 85.57 82.73 84.10 73.69
deb-Opus-MT 80.19 84.85 82.13 83.45 73.86
m2m100 81.27 85.89 84.09 84.96 74.10
ru Opus-MT 82.53 86.20 84.99 85.58 76.86
deb-Opus-MT 82.76 86.46 85.27 85.85 77.02
m2m100 81.71 86.39 84.06 85.21 75.21
Table 11: Comparison of comet Scores Overall, in Pro-/Anti-Stereotypical and Ambiguous Cases Across Models, on the 100 manually translated sentences.
Appendix I Log Probability and Surprisal Scores
Table 12 presents the Log probability and surprisal scores, as well as their relative differences between the Correct and Incorrect translations of the Unambiguous instances in WinoMT.
Language Model LogProb (Correct) LogProb (Incorrect) s3e $I$ (Correct) s3e $I$ (Incorrect) se $I$ (Correct) se $I$ (Incorrect) ge $I$ (Correct) ge $I$ (Incorrect) ES Opus-MT -149.7 -149.78 7.83 8.88 0.3 0.35 0.33 0.3 deb-Opus-MT -149.19 -149.01 8.08 9.16 0.29 0.31 0.35 0.33 m2m100 -226.61 -227.29 23.61 26.03 0.41 0.4 0.42 0.43 FR Opus-MT -197.1 -195.11 9.18 9.89 0.73 0.72 0.24 0.29 deb-Opus-MT -196.98 -195.09 9.18 9.85 0.48 0.52 0.26 0.33 m2m100 -283.91 -281.71 186.42 194.52 0.49 0.4 0.43 0.48 UK Opus-MT -161.98 -161.14 147.6 152.15 0.6 0.54 0.22 0.22 Opus-MT-debiased -161.46 -160.68 150.52 153.76 0.49 0.47 0.23 0.23 m2m100 -241.0 -241.49 204.72 211.9 0.28 0.24 0.23 0.25 RU Opus-MT -170.72 -170.9 32.14 33.15 0.38 0.43 0.08 0.19 deb-Opus-MT -170.58 -170.75 32.31 33.27 0.32 0.37 0.06 0.17 m2m100 -220.25 -220.78 218.11 219.06 0.45 0.4 0.16 0.3
Table 12: Log Probability and Surprisal Measures across Models and Languages
Appendix J Rankings by Different Metrics
Table 13 presents the rankings of models according to various metrics employed in this study.
comet Unambiguous comet All Gender Acc Delta S Delta H
deb-Opus-MT-RU Opus-MT-ES m2m100-ES m2m100-ES m2m100-RU
Opus-MT-RU deb-Opus-MT-ES deb-Opus-MT-ES deb-Opus-MT-ES m2m100-UK
m2m100-RU Opus-MT-FR Opus-MT-ES Opus-MT-ES Opus-MT-UK
m2m100-UK deb-Opus-MT-FR deb-Opus-MT-FR deb-Opus-MT-FR deb-Opus-MT-UK
Opus-MT-ES Opus-MT-RU Opus-MT-FR m2m100-FR Opus-MT-RU
deb-Opus-MT-ES deb-Opus-MT-RU m2m100-FR Opus-MT-FR deb-Opus-MT-RU
Opus-MT-UK m2m100-FR Opus-MT-RU deb-Opus-MT-RU deb-Opus-MT-ES
deb-Opus-MT-UK m2m100-ES m2m100-RU m2m100-RU Opus-MT-ES
Opus-MT-ES Opus-MT-UK deb-Opus-MT-RU deb-Opus-MT-UK deb-Opus-MT-FR
Opus-MT-FR deb-Opus-MT-UK m2m100-UK Opus-MT-UK m2m100-FR
deb-Opus-MT-FR m2m100-RU deb-Opus-MT-UK m2m100-UK m2m100-ES
m2m100-FR m2m100-UK Opus-MT-UK Opus-MT-RU Opus-MT-FR
Table 13: Model rankings across five evaluation metrics.
Appendix K Rank Correlation
Table 14 presents the correlation scores between $\Delta I$ and $\Delta$ Log Probabilies on the one hand, and gender accuracy scores on the other.
Correlation Metric Statistic p-value
Spearman $\Delta I$ (s3e) -0.78 0.00
$\Delta I$ (se) -0.37 0.24
$\Delta I$ (ge) 0.27 0.00
ΔLog $\Delta\text{Log prob}$ 0.11 0.73
Kendall $\Delta I$ (s3e) -0.58 0.01
$\Delta I$ (se) -0.27 0.25
$\Delta I$ (ge) 0.23 0.00
ΔLog $\Delta\text{Log prob}$ 0.09 0.74
Table 14: Spearman and Kendall correlations between $\Delta I$ under different uncertainty metrics and Log Probabilities on the one hand, and gender accuracy on the other. Statistically significant correlations (p $<$ 0.05) are in bold.
Appendix L Entropy Scores
Table 15 presents the $\mathcal{H}$ scores and their relative differences between the unambiguous and ambiguous settings for different UQ metrics used in this study.
Language Model s3e se ge Unamb. Amb. $\Delta\mathcal{H}$ Unamb. Amb. $\Delta\mathcal{H}$ Unamb. Amb. $\Delta\mathcal{H}$ ES Opus-MT 1.23 1.12 0.09 0.33 0.22 0.33 0.21 0.13 0.38 deb-Opus-MT 0.97 0.89 0.08 0.41 0.25 0.39 0.24 0.16 0.33 m2m100 1.79 1.45 0.19 0.46 0.17 0.63 0.25 0.09 0.64 FR Opus-MT 1.79 1.43 0.20 0.57 0.40 0.30 0.23 0.08 0.65 deb-Opus-MT 1.21 1.08 0.11 0.64 0.50 0.22 0.23 0.09 0.61 m2m100 3.22 2.78 0.14 0.56 0.29 0.48 0.25 0.17 0.32 UK Opus-MT 1.96 2.16 -0.10 0.40 0.39 0.03 0.20 0.14 0.30 deb-Opus-MT 1.98 2.15 -0.09 0.44 0.41 0.07 0.22 0.15 0.32 m2m100 2.05 2.28 -0.11 0.37 0.41 -0.11 0.18 0.16 0.11 RU Opus-MT 1.56 1.68 -0.08 0.38 0.32 0.16 0.12 0.03 0.75 deb-Opus-MT 1.05 0.97 0.08 0.43 0.42 0.02 0.12 0.06 0.50 m2m100 1.83 2.29 -0.25 0.50 0.29 0.42 0.15 0.10 0.33
Table 15: $\mathcal{H}$ scores across models and languages, with relative differences ($\Delta\mathcal{H}$) between unambiguous and ambiguous conditions.